# Matching-Based Policy Learning

### v1

@@ 2.3 Bias-Corrected Matching-Based Advantage Function
$\tilde{Y}_{i}(1)=\left\{\begin{aligned} &\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}\left(Y_{j}+\hat{\mu}(X_{i},1)-\hat{\mu}(X_{j},1)\right)&\text{ if }W_{i}=0,\\ &Y_{i}&\text{ if }W_{i}=1.\end{aligned}\right.$
Remarkably, they established the $\sqrt{n}$-asymptotic normality of the corresponding matching estimator $\hat{\tau}_{bc}=n^{-1}\sum_{i=1}^{n}\left(\tilde{Y}_{i}(1)-\tilde{Y}_{i}(1)\right)$ for $\tau$ under regularity conditions. Motivated by Abadie and Imbens (2011), we develop an improved matching-based advantage function:
$\hat{A}^{bc}_{match}(\pi)=\frac{1}{n}\sum^{n}_{i=1}\left(2\pi(X_{i})-1\right)\left(\tilde{Y}_{i}(1)-\tilde{Y}_{i}(0)\right).$ (3)
After basic manipulations, it follows that $\hat{A}^{bc}_{match}(\pi)=\hat{A}_{match}(\pi)-\hat{B}_{M}(\pi)$, where $\hat{B}_{M}(\pi)$ serves as an estimator of $B_{M}(\pi)$ given by
$\hat{B}_{M}(\pi)=\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left(2\pi(X_{i})-1\right)\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}\left(\hat{\mu}(X_{i},1-W_{i})-\hat{\mu}(X_{j},1-W_{i})\right).$
We proceed to establish the $\sqrt{n}$-asymptotic normality of $\hat{A}^{bc}_{match}(\pi)$ in Corollary 2, which implies that $\hat{A}^{bc}_{match}(\pi)-A(\pi)=O_{p}(1/\sqrt{n})$. Assumption 3 contains regularity conditions for the power series estimator $\hat{\mu}(x,w)$ to ensure its fast convergence rate and thus eliminate the conditional bias. Detailed discussions can be found in Abadie and Imbens (2011) and Otsu and Rai (2017).
@@
(iii) For each $\lambda$ and $w=0,1$, the derivative $\partial^{\|\lambda\|_{1}}\mu(x,w)/\partial x^{\lambda_{1}}_{1}\cdots\partial x^{\lambda_{p}}_{p}$ exists and its norm is bounded by $C^{\|\lambda\|_{1}}$ for some positive constant $C$.
Corollary 2.
Under Assumptions 1-3, then for any $\pi\in\Pi$,
$\left(V^{\tau(X)}(\pi)+V^{E}(\pi)\right)^{-1/2}\sqrt{n}\left(\hat{A}^{bc}_{match}(\pi)-A(\pi)\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N(0,1).$
2.4 From Estimating to Learning
@@
In the preceding discussion, we have introduced a bias-corrected matching-based estimator $\hat{A}^{bc}_{match}(\pi)$ of the advantage function $A(\pi)$ with a fixed policy $\pi$. We proceed to demonstrate the procedure of determining the optimal policy based on $\hat{A}^{bc}_{match}(\pi)$. Notice that $\hat{A}^{bc}_{match}(\pi)$ has the following expression:
$\hat{A}^{bc}_{match}(\pi)=\frac{1}{n}\sum^{n}_{i=1}\left(2\pi(X_{i})-1\right)\operatorname{sign}\left(\tilde{Y}_{i}(1)-\tilde{Y}_{i}(0)\right)\left|\tilde{Y}_{i}(1)-\tilde{Y}_{i}(0)\right|.$ (4)
@@
Therefore, the maximization for learning the optimal policy is equivalent to a weighted classification problem, where we can treat $\pi(\cdot)$ as the classifier to be trained, $\left\{\operatorname{sign}\left(\tilde{Y}_{i}(1)-\tilde{Y}_{i}(0)\right)\right\}^{n}_{i=1}$ as the labels, and $\left\{\left|\tilde{Y}_{i}(1)-\tilde{Y}_{i}(0)\right|\right\}^{n}_{i=1}$ as the weights.
@@
Given a set of $n$ training samples $\{X_{i},W_{i},Y_{i}\}_{i=1}^{n}$ and a restricted policy class $\Pi$, we summarize the implementation of MB-learning as follows:
@@
  • Impute the pairs of potential outcomes $\left\{Y_{i}(1),Y_{i}(0)\right\}^{n}_{i=1}$ using the bias-corrected matching estimators $\left\{\tilde{Y}_{i}(1),\tilde{Y}_{i}(0)\right\}^{n}_{i=1}$ defined in Section 2.3.
@@
  • Form the bias-corrected matching-based advantage function $\hat{A}^{bc}_{match}(\pi)$.
  • Implement the weighted classification (4) to learn the optimal policy $\hat{\pi}=\arg\max_{\pi\in\Pi}\hat{A}^{bc}_{match}(\pi)$.
@@
The idea of recasting the maximization of value function into a weighted classification framework has been previously discussed (e.g., Zhang et al., 2012a; Zhao et al., 2012; Athey and Wager, 2021). However, solving this maximization may be problematic since it is not a convex optimization problem. Following the argument in Athey and Wager (2021), we set the policy class $\Pi$ as a class of fixed-depth decision trees and implement the weighted classification problem using R package policytree (Sverdrup et al., 2020; R Core Team, 2023). This approach offers appealing interpretability and practical implementation for decision-makers. Some existing literature has also considered using convex surrogate loss function to replace the nonsmooth indicator function, thereby reformulating the policy learning problem into a convex optimization problem and implementing the optimization using off-the-shelf packages (Zhao et al., 2012; Liu et al., 2018).
3 Regret Bound
@@
In the previous section, we have investigated asymptotic behaviour of the proposed matching-based estimators for the advantage function $A(\pi)$ with a fixed $\pi$, focusing on the estimation properties. However, in statistical learning literature, one commonly used criterion to evaluate a policy is how effectively it minimizes the gap between the value function of the optimal policy and the policy we have learned (e.g., Zhao et al., 2012; Kitagawa and Tetenov, 2018; Athey and Wager, 2021; Zhou et al., 2023). This criterion is formalized through the concept of regret. Specifically, the utilitarian regret of a policy $\pi\in\Pi$ (Athey and Wager, 2021; Zhou et al., 2023) is defined as
@@
We proceed to investigate the convergence rate of $R(\hat{\pi})$, i.e., we will derive a high probability upper bound for the regret. Let $\text{VC}(\Pi)$ denote the Vapnik–Chervonenkis (VC) dimension (van der Vaart and Wellner, 2023) of the policy class $\Pi$. We assume that $\text{VC}(\Pi)$ is finite in Theorem 3 to control the complexity of $\Pi$. Suppose that we learn the optimal policy within a class of depth-$H$ decision trees in the implementation of MB-learning, then the corresponding VC dimension is $\tilde{O}\left(2^{H}\log p\right)$ (Athey and Wager, 2021). Here, the notation $\tilde{O}\left(f(n)\right)$ represents any factor bounded above by $g(f(n))f(n)$, where the function $g(\cdot)$ scales polylogarithmically with its argument. We also assume the countability of $\Pi$, which is a standard technical condition to simplify the measurability in the proof (e.g., Kitagawa and Tetenov, 2018). Let the notation $A\lesssim B$ denote that $A$ is less than $B$ up to a constant.
@@
Under Assumptions 1 and 3, suppose that the outcome $|Y|\leq L$ almost surely for some positive constant $L$, and the policy class $\Pi$ is countable with finite $\text{VC}(\Pi)$, then for any $\delta>0$ and any $q>1$, there exists an integer $N_{\delta,q}$ such that with at least probability $1-11\delta$ and for all $n>N_{\delta,q}$,
@@
It suffices to analyze the convergence rate of $A(\pi^{*})-A(\hat{\pi})$ in the proof, where we decompose it into several terms and derive the high probability upper bounds separately. We note that it could be challenging to derive the upper bound for
@@
due to the dependence among the terms $\left\{\sum_{j=1}^{n}\pi(X_{j})M_{ji}\right\}_{i=1}^{n}$, which is originated from the use of matching estimator. However, conditional on $\{X_{i},W_{i},\epsilon_{i}\}_{i=1}^{n}$, we can still utilize standard empirical process techniques including symmetrization (Boucheron et al., 2013) and demonstrate that the corresponding Rademacher process is sub-Gaussian. Consequently, we can leverage Dudley’s entropy integral bound (Vershynin, 2018) and the order of $\max_{i\leq n}K_{M}(i)$ to derive an upper bound for (5).
@@
Theorem 3 shows that the regret bound depends on the VC dimension $\text{VC}(\Pi)$, the number of matches $M$, and the quantity $V^{*}$. A better regret bound can be achieved if we decrease the complexity of the policy class or increase the size of matching. Analogous phenomenon concerning the relationship between the regret $R(\hat{\pi})$ and $M$ has been observed in Abadie and Imbens (2006, Theorem 5), where they stated that a larger $M$ can lead to a more efficient matching estimator of $\tau$. Recall that the unit-level term of $\bar{A}(\pi)$ is given by $\bar{A}_{i}(\pi)=(2\pi(X_{i})-1)\left(\mu(X_{i},1)-\mu(X_{i},0)\right)$, then $V^{*}$ can be reformulated as $\sup_{\pi_{1},\pi_{2}\in\Pi}E\left[\bar{A}_{i}(\pi_{1})-\bar{A}_{i}(\pi_{2})\right]^{2}$. Therefore, we can interpret $V^{*}$ as the worst-case discrepancy between the advantage functions of any two policies in the class $\Pi$. See Zhou et al. (2023) for detailed discussion.
From Theorem 3 we establish that the convergence rate of regret is
$R(\hat{\pi})=O_{p}\left(\sqrt{\frac{\text{VC}(\Pi)}{n^{1-1/q}}}\right),$
where $q>1$. We let $q$ go to $\infty$, then the convergence rate almost achieves $1/\sqrt{n}$. Though the convergence rate of MB-learning is slower than that of the method proposed by Athey and Wager (2021), which is exactly $1/\sqrt{n}$, our results in simulation studies and real data analysis demonstrate the superiority of the proposed method, especially in a relatively small sample size. Recall that M-learning proposed by Wu et al. (2020) also utilizes the matching for learning the optimal policy. However, they established the convergence rate based on the results of NN estimator (e.g., Devroye et al., 2013), where the number of matches grows with sample size. In contrast, the convergence rate of MB-learning is developed with a fixed number $M$. Additionally, Kitagawa and Tetenov (2018) showed that their IPWE-based method has a regret bound faster than $1/\sqrt{n}$ under the margin assumption, which is commonly used in classification literature (e.g., Tsybakov, 2004; Massart and Nédélec, 2006). Therefore, a faster convergence rate of MB-learning may be achieved under this assumption, but it is beyond the scope of this work.
@@ 4.1 Data Generating Process
We conduct extensive simulation studies to assess the finite sample performance of MB-learning. In each simulation setup, the covariates vector $X=(X_{1},X_{2},X_{3},X_{4})^{\top}$ is independently generated from $N(0,\mathbf{I}_{4})$. The binary treatment $W$ is generated from $\Pr(W=1|X)=\exp\left(l(X)\right)/\left[1+\exp\left(l(X)\right)\right]$, where we consider five scenarios of propensity score models based on different specifications of $l(X)$. Detailed information is presented in Table S1 of Supplementary Materials. The outcome variable is generated by $Y=m(X)+Wc(X)+Z,$ where $Z$ is the noisy variable independently generated from $N(0,1)$. We consider two scenarios of the main effect model $m(\cdot)$: Linear, $m(X)=1+2X_{1}-X_{2}+0.5X_{3}-1.5X_{4}$, and Nonlinear, $m(X)=4\sin(X_{1})+2.5\cos(X_{2})-X_{3}X_{4}$. The first main effect model is linear with its argument, while the second model exhibits highly nonlinearity. Additionally, two scenarios of the contrast function $c(\cdot)$ are also considered: Tree, $c(X)=2I\left\{X_{1}>0,X_{2}>0\right\}-1$, and Non-tree, $c(X)=2I\left\{2X_{2}-\exp(1+X_{1})+2>0\right\}-1$. We define the global optimal policy $\pi^{*}_{opt}$ as the optimal policy within all measurable functions from $\mathcal{X}$ to $\{0,1\}$ that maximizes the value function (Mo et al., 2021), i.e., $\pi^{*}_{opt}=\arg\max_{\pi}V(\pi)$. Note that the contrast function $c(X)$ determines the structure of the global optimal policy since $\pi^{*}_{opt}(X)=I\left\{c(X)>0\right\}$. Therefore, in the first scenario, we proceed to assess the performance of MB-learning when the restricted policy class $\Pi$ includes the global optimal policy $\pi^{*}_{opt}$. On the other hand, we investigate how the misspecification of $\Pi$ influences the proposed method in the second scenario. Similar simulation studies have also been conducted in the literature concerning tree-based policy learning approaches (e.g., Zhang et al., 2012a; Laber and Zhao, 2015; Athey and Wager, 2021).

### v2

@@ 2.3 Bias-Corrected Matching Estimator of the Advantage Function
$\tilde{Y}_{i}(1)=\left\{\begin{aligned} &\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}\left\{Y_{j}+\hat{\mu}(X_{i},1)-\hat{\mu}(X_{j},1)\right\},&\text{ if }W_{i}=0,\\ &Y_{i},&\text{ if }W_{i}=1.\end{aligned}\right.$
The bias-corrected matching estimator of the ATE $\tau$ is $\hat{\tau}_{bc}=n^{-1}\sum_{i=1}^{n}\{\tilde{Y}_{i}(1)-\tilde{Y}_{i}(1)\}$. Analogously, we develop a bias-corrected matching-based advantage function:
$\hat{A}^{bc}_{match}(\pi)=\frac{1}{n}\sum^{n}_{i=1}\left(2\pi(X_{i})-1\right)\left\{\tilde{Y}_{i}(1)-\tilde{Y}_{i}(0)\right\}.$ (3)
It follows that
$\hat{A}^{bc}_{match}(\pi)=\hat{A}_{match}(\pi)-\hat{B}_{M}(\pi),$
where the term
$\hat{B}_{M}(\pi)=\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left\{2\pi(X_{i})-1\right\}\left[\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}\left\{\hat{\mu}(X_{i},1-W_{i})-\hat{\mu}(X_{j},1-W_{i})\right\}\right]$
serves as an estimator of $B_{M}(\pi)$. The following assumption contains regularity conditions for the power series estimator $\hat{\mu}(x,w)$ to ensure its fast convergence rate and thus eliminate the conditional bias. Detailed discussions can be found in Abadie and Imbens (2011) and Otsu and Rai (2017).
@@
(iii) For each $\lambda$ and $w=0,1$, the derivative $\partial^{\|\lambda\|_{1}}\mu(x,w)/\partial x^{\lambda_{1}}_{1}\cdots\partial x^{\lambda_{p}}_{p}$ exists and its norm is upper bounded by $C^{\|\lambda\|_{1}}$ for some positive constant $C$.
We proceed to establish the asymptotic normality of $\hat{A}^{bc}_{match}(\pi)$ in the following theorem.
Theorem 2.
Under Assumptions 1-3, then for any $\pi\in\Pi$, as $n\rightarrow\infty$,
$\left\{V^{\tau(X)}(\pi)+V^{E}(\pi)\right\}^{-1/2}\sqrt{n}\left\{\hat{A}^{bc}_{match}(\pi)-A(\pi)\right\}\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,1).$
It follows immediately that $\hat{A}^{bc}_{match}(\pi)-A(\pi)=O_{p}(1/\sqrt{n})$, and thus the convergence rate of the bias-corrected matching estimator is faster when $p\geq 3$ than that in Corollary 1.
3 Policy Learning with Matching
@@
In the preceding discussion, we introduced matching estimators of the advantage function $A(\pi)$ when the policy $\pi$ is fixed. We now proceed to policy learning with matching. We propose to learn the optimal policy, denoted as $\hat{\pi}$, by optimizing the bias-corrected matching estimator of the advantage function $\hat{A}^{bc}_{match}(\pi)$ within a pre-determined policy class $\Pi$:
$\hat{\pi}=\arg\max_{\pi\in\Pi}\hat{A}^{bc}_{match}(\pi).$ (4)
We now proceed to describe the algorithm to solve the optimization problem (4). Furthermore, we derive a non-asymptotic regret bound for the learned policy $\hat{\pi}$.
3.1 Implementation Detail
Notice that $\hat{A}^{bc}_{match}(\pi)$ has the following expression:
$\hat{A}^{bc}_{match}(\pi)=\frac{1}{n}\sum^{n}_{i=1}\left(2\pi(X_{i})-1\right)\operatorname{sign}\left\{\tilde{Y}_{i}(1)-\tilde{Y}_{i}(0)\right\}\left|\tilde{Y}_{i}(1)-\tilde{Y}_{i}(0)\right|.$ (5)
@@ 3.1 Implementation Detail
Therefore, the problem of learning the optimal policy in (4) using the expression (5) is equivalent to a weighted classification problem, where we treat $\pi(\cdot)$ as the classifier to be trained, $\{\operatorname{sign}\{\tilde{Y}_{i}(1)-\tilde{Y}_{i}(0)\}\}^{n}_{i=1}$ as the labels, and $\{|\tilde{Y}_{i}(1)-\tilde{Y}_{i}(0)|\}^{n}_{i=1}$ as the weights.
@@
Given a set of $n$ training samples $\{X_{i},W_{i},Y_{i}\}_{i=1}^{n}$ and a restricted policy class $\Pi$, we describe the implementation of MB-learning as follows:
  1. 1.
@@
    Impute the pairs of potential outcomes $\{Y_{i}(1),Y_{i}(0)\}^{n}_{i=1}$ using the bias-corrected matching estimators $\{\tilde{Y}_{i}(1),\tilde{Y}_{i}(0)\}^{n}_{i=1}$ defined in Section 2.3.
  2. 2.
@@
    Construct the bias-corrected matching-based advantage function $\hat{A}^{bc}_{match}(\pi)$ in equation (5).
  3. 3.
    Solve the optimization problem (4) from the weighted classification perspective.
@@
The idea of recasting the optimization problem of the value function into a weighted classification framework has been previously discussed in the literature (e.g., Zhang et al., 2012a; Zhao et al., 2012; Athey and Wager, 2021). Directly solving this optimization problem in Step 3 is difficult since the estimated advantage function $\hat{A}^{bc}_{match}(\pi)$ is not convex. In this paper, we employ the policy tree algorithm in Athey and Wager (2021), where we restrict the policy class $\Pi$ to be a class of fixed-depth decision trees and solve the weighted classification problem using the R package policytree (Sverdrup et al., 2020). Due to its tree structure, the learned policy is easy to understand and interpret for decision-makers.
Alternatively, one may use some convex surrogate loss function to replace the nonsmooth indicator function in equation (5), thereby reformulating the policy learning problem into a convex optimization problem which can be solved using off-the-shelf packages (Zhao et al., 2012; Liu et al., 2018). This is beyond the scope of this work.
3.2 Regret Bound
@@
In the previous section, we proved the asymptotic behaviour of the proposed matching-based estimators for the advantage function $A(\pi)$ for a fixed policy $\pi$. In this subsection, we evaluate the performance of the learned policy compared to the underlying optimal policy $\pi^{*}$ using the criterion of regret. The utilitarian regret of a policy $\pi\in\Pi$ (Athey and Wager, 2021; Zhou et al., 2023) is defined to be
@@
We proceed to investigate the convergence rate of the regret $R(\hat{\pi})$ with the learned matching-based policy $\hat{\pi}$ obtained from the optimization problem (4). We aim to derive a non-asymptotic upper bound for $R(\hat{\pi})$ with high probability.
For a given policy class $\Pi$, let $\text{VC}(\Pi)$ be its Vapnik-Chervonenkis (VC) dimension (van der Vaart and Wellner, 2023). In the following theorem, We assume that $\text{VC}(\Pi)$ is finite to control the complexity of $\Pi$. Suppose that we learn the optimal policy by MB-learning within a class of depth-$H$ decision trees, then the corresponding VC dimension is $\tilde{O}\left(2^{H}\log p\right)$ (Athey and Wager, 2021). Here, the notation $\tilde{O}\left(f(n)\right)$ represents any factor upper bounded by $g(f(n))f(n)$, where the function $g(\cdot)$ scales polylogarithmically with its argument. We also assume that $\Pi$ is countable, which is a standard technical condition to simplify the measurability (e.g., Kitagawa and Tetenov, 2018). We write the notation $A\lesssim B$ to indicate that $A$ is less than $B$ up to a constant.
@@
Under Assumptions 1-3, suppose that the outcome $|Y|\leq L$ almost surely for some positive constant $L$, the policy class $\Pi$ is countable, and $\text{VC}(\Pi)$ is finite, then for any $\delta>0$ and any $q>1$, there exists an integer $N_{\delta,q}$ such that for all $n>N_{\delta,q}$, the regret $R(\hat{\pi})$ is upper bounded as follows with probability at least $1-11\delta$:
@@
It is equivalent to analyze the bound for $A(\pi^{*})-A(\hat{\pi})$ in the proof. We decompose it into several terms, for which we derive high-probability upper bounds, respectively. We remark that it is challenging to derive the upper bound for the following term:
$E\left[\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left\{\sum_{j=1}^{n}\pi(X_{j})M_{ji}\right\}\epsilon_{i}\right|\right],$ (6)
@@
because that matching induces pairwise correlations between the items $\{\sum_{j=1}^{n}\pi(X_{j})M_{ji}\}_{i=1}^{n}$. However, conditional on $\{X_{i},W_{i},\epsilon_{i}\}_{i=1}^{n}$, it is still eligible to utilize standard empirical process techniques including symmetrization (Boucheron et al., 2013) to show that the corresponding Rademacher process is sub-Gaussian. Consequently, we leverage Dudley’s entropy integral bound (Vershynin, 2018) and the order of $\max_{i\leq n}K_{M}(i)$ to derive an upper bound for the term (6).
@@
Theorem 3 shows that the regret bound depends on the VC dimension $\text{VC}(\Pi)$, the number of matches $M$, and the quantity $V^{*}$. A better regret bound can be achieved if we decrease the complexity of the policy class or increase the number of matching. Similar discussion can be found in Abadie and Imbens (2006, Theorem 5), where they stated that a larger $M$ generally leads to a more efficient matching estimator of the ATE. Recalling that the unit-level term of $\bar{A}(\pi)$ is given by $\bar{A}_{i}(\pi)=\{2\pi(X_{i})-1\}\{\mu(X_{i},1)-\mu(X_{i},0)\}$, the term $V^{*}$ can be reformulated as $\sup_{\pi_{1},\pi_{2}\in\Pi}E\left[\bar{A}_{i}(\pi_{1})-\bar{A}_{i}(\pi_{2})\right]^{2}$. Therefore, we can interpret $V^{*}$ as the worst-case discrepancy between the advantage functions of any two policies in the class $\Pi$. See Zhou et al. (2023) for related discussions.
A direct consequence of Theorem 3 is the convergence rate of the regret $R(\hat{\pi})$:
Corollary 2.
Under the assumptions in Theorem 3, for any $q>1$, it holds that
$R(\hat{\pi})=O_{p}\left(\sqrt{\frac{\text{VC}(\Pi)}{n^{1-1/q}}}\right).$
Athey and Wager (2021) proved that the regret of AIPW-based policy learning with cross-fitting is rate-optimal with the order of $\sqrt{\text{VC}(\Pi)/n}$. Corollary 2 suggests that the regret of the proposed MB-learning has slower convergence rate for any $q>1$. Nevertheless, with $q$ going to $\infty$, the convergence rate is almost of order $\sqrt{\text{VC}(\Pi)/n}$. That is, MB-learning is almost rate-optimal.
Although MB-learning is not rate-optimal, the simulation studies and real data analysis demonstrate the superiority of the proposed method compared to AIPW-based policy learning, especially when the sample size is not large or when the propensity score has extreme values.
Kitagawa and Tetenov (2018) showed that their IPW-based method has a regret bound faster than the order of $1/\sqrt{n}$ under the margin assumption, which is commonly used in classification literature (e.g., Tsybakov, 2004; Massart and Nédélec, 2006). A faster convergence rate of the regret of MB-learning may be feasible under additional assumptions, but it is beyond the scope of this work.
@@ 4.1 Data Generating Process
We conduct extensive simulation studies to assess the finite sample performance of MB-learning. In each simulation setup, the covariate vector $X=(X_{1},X_{2},X_{3},X_{4})^{\top}$ is independently generated from $N(0,\mathbf{I}_{4})$. The binary treatment $W$ is generated from the propensity score $\Pr(W=1|X)=\exp\left(l(X)\right)/\left[1+\exp\left(l(X)\right)\right]$, where we consider five different specifications of $l(X)$. Detailed information is presented in Table S1 of Supplementary Material. The outcome variable is generated by $Y=m(X)+Wc(X)+e,$ where $e$ is a noise term generated from $N(0,1)$. We consider two scenarios of the main effect model $m(\cdot)$: (i). Linear: $m(X)=1+2X_{1}-X_{2}+0.5X_{3}-1.5X_{4}$; (ii). Nonlinear: $m(X)=4\sin(X_{1})+2.5\cos(X_{2})-X_{3}X_{4}$. Moreover, We consider two scenarios of the contrast function $c(\cdot)$: (i). Tree: $c(X)=2I\left\{X_{1}>0,X_{2}>0\right\}-1$; (ii). Non-tree: $c(X)=2I\left\{2X_{2}-\exp(1+X_{1})+2>0\right\}-1$. We define the global optimal policy $\pi^{*}_{opt}$ as the policy within all measurable functions from $\mathcal{X}$ to $\{0,1\}$ that maximizes the value function (Mo et al., 2021), i.e., $\pi^{*}_{opt}=\arg\max_{\pi}V(\pi)$. Note that the contrast function $c(X)$ determines the structure of the global optimal policy since $\pi^{*}_{opt}(X)=I\left\{c(X)>0\right\}$. Therefore, when the contrast function has the tree structure, the restricted policy class $\Pi$ includes the global optimal policy $\pi^{*}_{opt}$ in MB-learning. When the contrast function is non-tree, the policy class $\Pi$ is misspecified, and we investigate how the misspecification of $\Pi$ influences the proposed method. Similar simulation studies were conducted concerning tree-based methods in the literature (e.g., Zhang et al., 2012a; Laber and Zhao, 2015; Athey and Wager, 2021).